A chemist has three different acid solutions. The first acid solution contains 25% acid, the second contains 45% and the third contains 85% . He wants to use all three solutions to obtain a mixture of 50 liters containing 65% acid, using 3 times as much of the 85% solution as the 45% solution. How many liters of each solution should be used?
Answer by htmentor(1343)

Let x = the amount of the 45% acid solution
Then the amount of the 85% acid solution = 3x
Since the total volume of the mixture is 50 l, the amount of the 25% solution is 50 - 3x - x = 50 - 4x
The equation for the total amount of acid in the mixture is
0.65*50 = 0.45x + 0.85*3x + 0.25(50 - 4x)
Simplify and solve for x:
0.45x + 2.55x + 12.5 - x = 32.5
2x = 20
x = 10
So there 10 l of 45% solution, 30 l of 85% solution and 10 l of 25% solution